OpenStudy (anonymous):

Lucy is financing $370,000 to purchase a house. How much money will she save over the life of a 30-year, fixed-rate loan by buying 3 points with a rate of 6.175% instead of not buying points with a rate of 6.55%? $21,548.40 $14,364.84 $3,264.84 $32,648.40

OpenStudy (anonymous):

@FriedRice

OpenStudy (anonymous):

1) 3 points=3% x 370000 = 11,000 so PV=359000 Payment =PV(i/n)/[1 - [[1 /1+ i/n]]^t*n] P=359000(.06175/12)/[1-[[1/1+0.06175/1… 359000(.00515)/[1-[[.99488]^360]] 1848.85/[1-.15759] pmnt=2,194.72 you can do it for the 6.55% and this should work for most of the other questions

OpenStudy (anonymous):

Like what do I do from here to find the answer???

OpenStudy (anonymous):

I = PRT I = Interest P = Percent R = Rate T = Time

OpenStudy (anonymous):

P means The amount

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So can you help me solve it

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what will be the interest

OpenStudy (anonymous):

6.175%?

OpenStudy (anonymous):

that's the interest rate

OpenStudy (anonymous):

ok we got this part I = P(6.175)(T)

OpenStudy (anonymous):

ok now what is the amount?

OpenStudy (anonymous):

$370,000 is the amount we got I = (370000)(6.175)(T)

OpenStudy (anonymous):

and the time is 30 years so I = (370000)(6.175)(30) so what would the Interest be or I?

OpenStudy (anonymous):

68542500

OpenStudy (anonymous):

ok then devide what we got last time 2,194.72

OpenStudy (anonymous):

so its closest to D that must be your answer

OpenStudy (anonymous):

thank you so much can you help with a few more please??

OpenStudy (anonymous):

I g2g

OpenStudy (anonymous):

sorry ill call hero or someone else

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

Ok I am going to post them hopefully they help...

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

if you have another question, you should make another post.